If the demand function for rice is as follows

Qd = -200P + 5000

Find the consumer surplus when the price is fixed at

a) GHc 15

b) GHc 10
1
Expert's answer
2020-10-19T13:36:52-0400
"Q_1=-200\times15+5000=2000"


"Q_2=-200 \times 10+5000=3000"